OpenStudy (anonymous):

MEDAL 7 : 49 What it the above ratio in lowest terms

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

@paki

OpenStudy (kweenlexis):

7 / 49 = 7 / 72 = (7 : 7) / (72 : 7) = (7 : 7) / (49 : 7) = 1 / 7